A car with a mass of 900.0 kg is moving at 26.0 m/s [E] when it crashes into a truck with a mass of 1500.0 kg moving at 30.0 m/s [W]. After the 0.4 s collision, both vehicles are coupled together and have the same velocity. Find the final velocity and average net force of each vehicle after the collision.


1
Expert's answer
2021-02-22T10:25:38-0500

(a) Let the east be the positive direction. Then, according to the law of conservation of momentum, we have:


"m_1v_1-m_2v_2=(m_1+m_2)v_f,""v_f=\dfrac{m_1v_1-m_2v_2}{(m_1+m_2)},""v_f=\dfrac{900\ kg\cdot26\ \dfrac{m}{s}-1500\ kg\cdot30\ \dfrac{m}{s}}{(900\ kg+1500\ kg)}=-9\ \dfrac{m}{s}."

The sign minus means that the couple of the car and truck moves to the west.

(b) Let's find the average net force of each vehicle after the collision:


"F_{avg}\Delta t=m|(v_f-v_i)|,""F_{avg}=\dfrac{m|(v_f-v_i)|}{\Delta t},""F_{avg, car}=\dfrac{900\ kg\cdot|(-9\ \dfrac{m}{s}-26\ \dfrac{m}{s})|}{0.4\ s}=78750\ N,""F_{avg, truck}=\dfrac{1500\ kg\cdot|(-9\ \dfrac{m}{s}-30\ \dfrac{m}{s})|}{0.4\ s}=146250\ N."